OpenStudy (anonymous):

SOLVE: please please t+7= radical 2t+13 end radical

OpenStudy (anonymous):

no one likes you everybody hates you obvi

OpenStudy (anonymous):

you know what Kirstin... HUSH CHILD!

OpenStudy (anonymous):

t+7=sqrt(2t+13) now square both sides (t+7)^2=2t+13 expand the left side of the equation t^2+14t+49=2t+13 t^2+12t+36=0 (t+6)(t+6)=0 so t=-6

OpenStudy (anonymous):

THANK YOU SO MUCH!!